Susan and jayne have part-time jobs selling tickets at the theatre. This week, Jayne s paycheck was twice the amount of susan s paycheck. They both were paid 7.50 an hour, and jaynes paycheck was $60. How many hours did susan work this week? (susan earns half the amount jayne earns)



if possible i was wondering if you could do this problem step by step so i will know how to do it and please respond back too it as quickly as possible?

Answer by checkley77(12844)

J=60/7.50=8 HOURS FOR jAYNE.
S=30/7.50=4 HOURS FOR SUSAN